OpenStudy (anonymous):

how do you divide 6 5/8 by 6 1/4

OpenStudy (anonymous):

you need to start by making each into an improper fraction... like 1 1/2 is the same as 3/2

OpenStudy (anonymous):

How many 8ths is 6 5/8 ?

OpenStudy (anonymous):

Each of the 6 whole numbers is eight 8ths, right? (8/8) = 1 So 6 is equivalent to 6 x (8/8) = 48 / 8 It is like 6 pies sliced into 8 slices each... there are 48 total slices

OpenStudy (anonymous):

So make 6 5/8 into (48/8) + (5/8) or = (53/8) and 6 1/4 into (24/4) + (1/4) = (25/4)

OpenStudy (anonymous):

then you have two fractions.. one divided by the other is the same as the first TIMES the inverse of the 2nd... and the inverse is just flipping the fraction over... So it's (53/8) X (4/25)

OpenStudy (anonymous):

That should be enough... :) hope this helps!!